Ask your own question, for FREE!
Mathematics 12 Online
OpenStudy (anonymous):

1) A die roll twice .Find probability P( two even number ) 2) Three cards are drawn from a standard deck of cards .Find each probability P(3 hearts) if replacement occurs P(3 hearts) i no replacement occurs

OpenStudy (anonymous):

1) 6 sided dice are 50% even numbers, so your probability is .5*.5=.25 2) 52 cards in a standard deck of cards, 13 are hearts with replacement, you get 13/52*13/52*13/52=1/4*1/4*1/4=1/64 without replacement you get 13/52*12/51*11/50=11/850

OpenStudy (anonymous):

looks good to me!

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
2 days ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
2 days ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!